A submarine is traveling at 9 m.p.h. and is descending at an angle of depression of 5 degrees, how many minutes does it take for the submarine to reach 120 feet?

Answer by lwsshak3(11628)

A submarine is traveling at 9 m.p.h. and is descending at an angle of depression of 5 degrees, how many minutes does it take for the submarine to reach 120 feet?
..
let x=minutes it takes submarine to reach a depth of 120 ft.
What we have here is a right triangle with a leg of 120 ft opposite the 5 deg angle and h=hypotenuse.
sin 5 deg = 120/h
h=120/sin 5 deg=1376.85 ft
travel time = distance/speed
x=(1376.85 ft/9 mi/hr)*(mi/5280 ft)*(60min/hr)
=1376.85*60/9*5280
=1.74 minutes
ans:
It takes the submarine 1.74 minutesto reach a depth of 120 ft.